OpenStudy (anonymous):

PLEASE HELP!!! How do you factor 4x^2 − 6x − 4?!?!?!?!?!?!

OpenStudy (trojanpoem):

use the quadratic formula -b + or - sqrt(b^2 - 4ac) / 2a a is the coefficient of x^2 b is the coefficient of x c is the coefficient of x^0 (const)

OpenStudy (anonymous):

can you show me how to do that?

jhonyy9 (jhonyy9):

so look it here please what terms are common 2x*2x -2x-2x-2x-2-2 ?

OpenStudy (anonymous):

2

OpenStudy (anonymous):

and 2x

jhonyy9 (jhonyy9):

yes so than factorize it out from every terms what will get ? first just 2

jhonyy9 (jhonyy9):

courage this is easy

OpenStudy (anonymous):

2(2x^2 - 3x - 2)

jhonyy9 (jhonyy9):

nice so now you need working with terms inside parentheses 2x^2 -3x-2 = 2x*x -x-2x-2 = grouping them from first 2 terms what can you factoriz out and from 3rd and 4th ?

OpenStudy (anonymous):

2

jhonyy9 (jhonyy9):

so 2 just from 3rd and 4th but from first and second ?

jhonyy9 (jhonyy9):

so for 2x^2 -3x -2 make it equal zero and using discriminant method can you get the roots of this quadratic ?

OpenStudy (anonymous):

what?

jhonyy9 (jhonyy9):

what is formula for discriminant ?

OpenStudy (anonymous):

the quadratic formula

jhonyy9 (jhonyy9):

D=b^2-4ac can you calculi it ?

OpenStudy (anonymous):

-28

jhonyy9 (jhonyy9):

how you have got it ?

jhonyy9 (jhonyy9):

b^2 = ?

OpenStudy (anonymous):

-6

OpenStudy (anonymous):

oh sorry 36

jhonyy9 (jhonyy9):

no you work now just with quadratic inside paretheses

jhonyy9 (jhonyy9):

parentheses - sorry

jhonyy9 (jhonyy9):

2x^2 -3x -2 =0 D=?

OpenStudy (anonymous):

I'm sorry, I don't understand

jhonyy9 (jhonyy9):

so first in case of initial quadratic you have factorized out 2 - yes ?

OpenStudy (anonymous):

right

OpenStudy (anonymous):

oh, I get it

jhonyy9 (jhonyy9):

than calculi it please

OpenStudy (anonymous):

okay hold on

jhonyy9 (jhonyy9):

can you solve it sure till end ?

OpenStudy (anonymous):

is it 25

OpenStudy (anonymous):

?

jhonyy9 (jhonyy9):

yes and x_1,2 = ?

OpenStudy (anonymous):

what does x_1,2 mean?

jhonyy9 (jhonyy9):

the roots of this quadratic inside parentheses

OpenStudy (anonymous):

oh

OpenStudy (anonymous):

they're -1/2 and 2

OpenStudy (anonymous):

right?

jhonyy9 (jhonyy9):

how you have got these roots ?

jhonyy9 (jhonyy9):

-b +/- sqrt(D) x_1,2 = -------------- 2a do you know this formula ?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

wait, is that the quadratic formula?

jhonyy9 (jhonyy9):

yes

jhonyy9 (jhonyy9):

so using this formula what roots will get ?

OpenStudy (anonymous):

um, hold on

jhonyy9 (jhonyy9):

2x^2 -3x -2 =0 D=25 3 +/- 5 x_1,2 = --------- = 2 and -1/2 4 right ?

jhonyy9 (jhonyy9):

so this mean that your results was right but how you have got them ?

OpenStudy (anonymous):

yes I got -1/2 and 2

jhonyy9 (jhonyy9):

so now using these roots you can rewriting the quadratic factorized -

jhonyy9 (jhonyy9):

how will be ?

OpenStudy (anonymous):

um, I'm not sure

jhonyy9 (jhonyy9):

why ? try it please

jhonyy9 (jhonyy9):

us formula

OpenStudy (anonymous):

what do you mean by rewriting the quadratic formula?

jhonyy9 (jhonyy9):

rewriting the quadratic using formula (x-x_1)(x-x_2) do you know this ?

OpenStudy (anonymous):

oh, so (x - 2)(x + 1/2) ?

jhonyy9 (jhonyy9):

so for you can being sure that is right you can check it - yes ?

OpenStudy (anonymous):

how do you check it?

jhonyy9 (jhonyy9):

multiplie term by term and look than you will get the necessary quadratic

OpenStudy (anonymous):

okay I got x^2 - (3x/2) - 1, but that's not the original

jhonyy9 (jhonyy9):

(x-2)(x+ 1/2) = x^2 +x/2 -2x -2/2 = yes for you can eliminate the denominator what you need to do ?

jhonyy9 (jhonyy9):

2x^2 +x -4x -2 -------- + -------- = ? 2 2

OpenStudy (anonymous):

(2x + 1)(x - 2)/2

jhonyy9 (jhonyy9):

in numerator you get 2x^2 -3x -2 the quadratic from what you have started your calculi than this mean that this quadratic is factorized completly will be (x-2)(x + 1/2) but dont forget that in the first step you have factorized the 2 so than what will be your exercise solved right ?

OpenStudy (anonymous):

2(x - 2)(x + 1/2) ?

jhonyy9 (jhonyy9):

yes but than you dont believe it make the prove of making the calculi multiplid term by term and you need getting the original quadratic ok ?

OpenStudy (anonymous):

ok.

OpenStudy (anonymous):

thank you :)

jhonyy9 (jhonyy9):

was my pleasure - hope so much that now you understand it sure right how you need solving same exercises - yes ?

OpenStudy (anonymous):

yes I understand it now

jhonyy9 (jhonyy9):

ok. good luck bye